OpenStudy (anonymous):

Probability question: Out of 50 parts, 10 are defective. If you choose 1 at random, what is probability 1st choice is defective? What is probability 1st and 2nd choice defective? What is probability 2nd choice defective, given 1st choice defective? What is probability second choice defective?

OpenStudy (anonymous):

I would like to check my answers....

OpenStudy (anonymous):

Probability 1st choice defective = 10/50 = 1/5 Probability 1st and 2nd choices defective = (1/5)(9/49) Probability 2nd defective given first is defective = 9/49

OpenStudy (anonymous):

Am I on the right track??

OpenStudy (anonymous):

Oh yes! Good job!

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Yep! However, the problem has a bit of ambiguity. Do you take the defective part off once you select it or not? It's not your fault, it's the problem's fault :)

OpenStudy (anonymous):

It was specified that you take if off and don't replenish (; Sorry I should have specified.